OpenStudy (anonymous):

Find the inverse function f(x)=e^(2x-1)

hartnn (hartnn):

let y = e^(2x-1) now express x in terms of y. can u? then just replace y with x.

OpenStudy (anonymous):

how do you express x in terms of y on this one

OpenStudy (lgbasallote):

@hartnn the math lover....are inverse functions expressed as y = ?

hartnn (hartnn):

take natural log on both sides. ln y = ln (e^(2x-1))

OpenStudy (anonymous):

now what hart

hartnn (hartnn):

u know the property of log, ln a^b = b ln a ??

OpenStudy (anonymous):

@lgbasallote this is how they teach us, to replace the f(x) with y then solve for x, then switch them when done

OpenStudy (lgbasallote):

this is why i hate math...too many rules

OpenStudy (anonymous):

im rusty with logs hartnn

hartnn (hartnn):

if u know then ln(e^(2x-1)) = (2x-1) ln e and ln of e =1 because the base of ln is also e.

hartnn (hartnn):

so u have ln y = 2x-1 now can u find x in terms of y ??

OpenStudy (anonymous):

i will explain let f(x)=y then, y= e^(2x-1), the most important thing is inverse function is symmetry to y=x so you can replace y to the x and x to the y so given funtion is arraged by x=e^(2y-1) by the rule y=e^x <=> lny=x lnx=2y-1 so answer is (lnx + 1)/2 = y

OpenStudy (anonymous):

perfect thanks guys

hartnn (hartnn):

ok, welcome :)

hartnn (hartnn):

ok, welcome :)